OpenStudy (waffelgirl12345):

A problem states: "There are 9 more children than parents in a room. There are 25 people in the room in all. How many children are there in the room?" What are the unknowns in this problem? Select each correct answer. the total number of parents in the room the total number of children in the room the total number of people in the room The total number of teachers in the room

OpenStudy (itz_sid):

\[C + P = 25\] \[P = (9+C)\] I believe

OpenStudy (sweetburger):

y + x =25, y= (9+x) let x represent parents let y represent children

OpenStudy (itz_sid):

Yes, then just solve for the children after substituting in the variables.

OpenStudy (itz_sid):

C = Children P =Parents \[C + P = 25\] \[P = 9 + C\] Then plus in your P, so you get \[C + (9+C) = 25\] Can you solve for C from there?

jimthompson5910 (jim_thompson5910):

@iTz_Sid you have the second equation a bit off. It should be C = P+9

jimthompson5910 (jim_thompson5910):

There are more children

OpenStudy (itz_sid):

@jim_thompson5910 Oh yea, sorry, Wooopsies.

OpenStudy (waffelgirl12345):

Thank u all for trying to help me...

OpenStudy (waffelgirl12345):

can anyone help me at all?

OpenStudy (waffelgirl12345):

plz someone help?????

OpenStudy (waffelgirl12345):

@helpppp

OpenStudy (itz_sid):

\[(P+9)+P = 25\] \[2P=16\] \[P=8\] Then plus P back in to get C because C = P+9 C = 8 + 9 C = 17

OpenStudy (mathmale):

A problem states: "There are 9 more children than parents in a room. There are 25 people in the room in all. How many children are there in the room?" What are the unknowns in this problem? Select each correct answer. the total number of parents in the room the total number of children in the room the total number of people in the room The total number of teachers in the room First and foremost, we need to choose variables to represent the unknown quantities. C is quite suitable to represent the number of students, P the number of parents. I'd choose N to represent the total number of people in the room. Note that we are told nothing about the number of teachers who are in the room. Note carefully: the goal of this problem is not so much to determine the exact numbers of people in each group (Children, Parents, Teachers, etc.), but rather to IDENTIFY THE UNKNOWNS. Do we know the total number of people in the room? Yes. We're told that it's 25. Thus, "total number of people in the room" is NOT an UNKNOWN here. Do we know the total number of Parents? No. Therefore, the "number of parents" is an unknown. Continue this pattern. Are there any other unknowns?

OpenStudy (waffelgirl12345):

that isnt 1 of my choices

OpenStudy (itz_sid):

Heh.... I prolly messed up.

OpenStudy (waffelgirl12345):

heh?? i dont get any of that!

OpenStudy (mathmale):

@waffelgirl12345 : Did you actually read my long previous message? Do you really know what the object of this problem is?

OpenStudy (mathmale):

Please review the problem statement. Find the part that tells you what you have to do. Explain in your own words what you are trying to find / identify.

jimthompson5910 (jim_thompson5910):

`There are 25 people in the room in all.` is this known or unknown?